# Ringdown signatures of Kerr black holes immersed in a magnetic field

### v2

@@ 3.1 Numerical QNM results
The equation can be solved using a root finding algorithm beginning with some initial guess $\omega^{0}_{{\ell mn}}$. To ensure convergence of the mode $\omega_{\ell mn}$ we will use the technique outlined in [60, 61, 62] to fine-tune the required number of terms $N$ needed in the continued fraction. In each step we begin with the starting value of $N$ and determine $\omega_{\ell mn}$, increasing $N$ by $\delta N$, until the relative difference between the last two adjacent values meets a specified precision threshold. That is, for a chosen set of parameters $\lambda$, if $\omega_{\ell mn}(N)$ denotes the QNM frequency determined with an $N$-term continued fraction, the iteration will stop when
@@
Earlier calculations in the literature for Kerr [56, 3] and Ernst [63, 64] black holes were used to validate our implementation of the continued fraction method. The non-Hermitian and thus inherently dissipative nature of the QNM (quasi)spectrum calculation requires the use of high machine precision throughout intermediate steps to avoid rounding errors, and in this work we have set the convergence criterion to $\epsilon=-7$.
We use two iterations to determine the QNM frequencies $\omega_{{\ell mn}}$ as a function of $\tilde{a}$ and $\tilde{B}$. The first iteration begins by setting $\ell,m,\text{ and }n$ and simultaneously fixing $\tilde{a}=0$. The parameter space for $\tilde{B}$ is then scanned incrementally through the range $(0,0.3)$. In this work we take the maximum value of the magnetic field to be $\tilde{B}=0.3$, given the constraints of the linearized model. Using the Schwarzschild QNM $\omega_{\ell 0n}=\omega^{0}_{\ell mn}$ as the initial starting guess in the root finding algorithm, the parameter space of $\tilde{B}$ is determined in steps of $\delta\tilde{B}=0.01$ up to the point $(\tilde{a},\tilde{B})=(0,0.3)$. A list of values $(0,\tilde{B},\omega_{\ell mn},N_{\mathrm{opt}})$ is constructed where $N_{\mathrm{opt}}$ is the optimal value of $N$ ensuring the convergence criterion defined in Eq. 34. The second iteration involves using this constructed list as the seed solution while scanning $\tilde{a}$. More precisely, for a given value of $\tilde{B}$, the list constructed in the first iteration is fed into the root finding algorithm while incrementally varying $\tilde{a}$, starting from $\tilde{a}=0.001$ which corresponds to a slowly rotating Kerr black hole up to $\tilde{a}=0.3$ in steps of $\delta\tilde{a}=0.01$. For each value of $\tilde{B}$, the initial starting guess is given by the result for $\omega^{0}_{{\ell mn}}=\omega_{\ell mn}(\tilde{a}=0,\tilde{B})$ and the $N_{\mathrm{opt}}$ value are used as combined starting points to speed up calculations in the second iteration while only the spin parameter is varied. This procedure results in a list of values $(\tilde{a},\tilde{B},\omega_{\ell mn})$ for the linearized Ernst-Wild black hole, exhibited graphically for several modes in Fig. 3.
A notable feature is that increasing $\tilde{B}$ leads to a reduction in the imaginary part of the QNM frequency in all cases. This reflects the tendency of the pure Melvin geometry to define box-like boundary conditions and exhibit normal as opposed to quasinormal modes. Another notable feature is the symmetry between positive and negative branches of modes. An analogous symmetry was observed by Leaver in [56] in the case of the Kerr geometry, and can be understood in a similar way. In particular, if $\omega_{\ell mn}$ is a QNM frequency with azimuthal index $m$, then $\omega_{\ell-mn}=-\omega_{\ell mn}^{*}$. Similarly, the linearized EW equations contain combinations of $m^{2}\tilde{B}^{2}$, and $m\tilde{a}\tilde{B}^{2}$ meaning that if $m\rightarrow-m$ then $m^{2}\tilde{B}^{2}$ should remain invariant while $m\tilde{a}\tilde{B}^{2}\rightarrow-m\tilde{a}\tilde{B}^{2}$. These symmetries are apparent on computing the positive and negative branch of the quasinormal mode frequencies, as shown in Fig. 3.
@@ 3.2 Numerical QNM interpolation
where $\omega_{\ell 0n}$ is the Schwarzchild QNM for a given  and $\ell\text{ and }n$ as listed in Table 1. We compute the coefficients $a_{\ell mn}$, $b_{\ell mn}$, and $c_{\ell mn}$ by taking numerical derivatives of the data obtained in the preceding section. In particular, $a_{\ell mn}$ describe the first derivative of the Kerr spectrum in the linearized regime of $\tilde{a}$, $b_{\ell mn}$ describe the first derivative of the Ernst spectrum in the linearized regime of $\tilde{B}^{2}$, and $c_{\ell mn}$ result from the numerical second derivative with respect to both $\tilde{a}$ and $\tilde{B}^{2}$. The results for these coefficients are tabulated in Tables 2 and 3, with those restricted to linearized Kerr agreeing with [54]. A comparison between the modes obtained directly from the Ernst geometry at ${\cal O}(\tilde{B}^{2})$ and the coefficients $b_{lmn}$ is shown in Fig. 4.
@@
1 $-1$ 0 -0.097660 -0.000336 0.268150 0.318134
$1$0.097660 0.000336 0.268150 -0.318132
2 $-2$ 0 -0.096759 -0.000143 0.432264 0.535567
$-1$0.096759 -0.000072 0.108133 0.132246
$1$0.096759 0.000072 0.108133 -0.132246
$2$0.096759 0.000143 0.432264 -0.535561
2 $-2$ 1 -0.295604 -0.004641 0.933283 1.131990
$-1$0.295604 -0.002321 0.233458 0.322617
$1$0.295604 0.002321 0.233458 -0.322418
$2$0.295604 0.004642 0.933283 -1.132600
3 $-3$ 0 -$0$0.000078 0.512260 0.653273
$-2$$0$0.000052 0.227786 0.263341
$-1$$0$0.000026 0.056964 0.080038
$1$0.096500 0.000026 0.056964 -0.080027
$2$0.096500 0.000052 0.227786 -0.263310
$3$0.096500 0.000078 0.512260 -0.653221
4 $-4$ 0 -0.096392 -0.000048 0.558610 0.726651
$-3$0.096392 -0.000036 0.314251 0.370947
$-2$0.096392 -0.000024 0.139677 0.164431
$-1$0.096392 -0.000012 0.034921 0.057358
$1$0.096392 0.000012 0.034921 -0.057357
$2$0.096392 0.000024 0.139677 -0.164431
$3$0.096392 0.000037 0.314251 -0.370944
$4$0.096392 0.000049 0.558610 -0.726644
Presently, there is some debate within the gravitational wave literature regarding which modes can be reliably extracted from the black hole ringdown. Indeed, it is clear that the process of ringdown following a physical binary merger is complex, and nonlinear effects may be as large as higher-order QNM contributions. In [15], it is argued that even in the ideal situation when a binary black hole system collides head-on, only the modes $\{(2,2,0),(3,2,0),(4,2,0)\}$ significantly contribute to the strain. Reliable extraction of the first overtone $(2,2,1)$ would then depend on a high-precision determination of the longer-lived fundamental modes of subdominant multipoles. In the non-ideal case when the binary does not collide head-on, it becomes significantly more difficult to extract overtones and the dominant modes excited could differ from $\{(2,2,0),(3,2,0),(4,2,0),(2,2,1)\}$. For this reason we include the fit coefficients for the fundamental modes from $1\leq\ell\leq 4$ as well as the first overtone for $\ell=2$ to cover the range expected to be relevant in a range of ringdown analyses.
@@ 4 Ringdown template and comparison to LVK data
In this section we utilize our QNM results for the linearized EW geometry to build and test a ringdown template against the existing LVK data archive of binary black hole mergers. This serves as a first step in modifying the usual Kerr template to model the post-merger phase of a binary black hole coalescence in the presence of an external asymptotically uniform magnetic field. We model our approach on [65], which considered the ringdown phase of black holes in the presence of a remnant $U(1)$ charge.
To investigate the magnetized Ernst-Wild hypothesis in the current LVK data archive, we make use of pyRing [66, 67, 68], a python software package specifically designed for the estimation of black hole ringdown parameters. pyRing implements a Bayesian analysis approach, described within the LVK analysis pipeline in Refs. [69, 70, 71, 72, 73, 74].
@@ 4.1 Constructing the ringdown template
To construct a waveform model for a black hole in the presence of an external magnetic field, we start from the standard Kerr template [65, 75, 76]:
@@
Here, $\tilde{\omega}_{s{\ell mn}}$ are the complex tensor ($s=-2$) ringdown frequencies of the Kerr black hole determined by the dimensionless final spin $\tilde{a}_{f}$ and the detector frame final mass given by $M_{f}$ defined in [77]. $M_{f}$ is related to the source frame mass by a factor of $(1+z)$, and $D_{L}$ is the luminosity distance. The $f$ subscript on parameters in this section indicates that these values refer to the final black hole of the merger, commonly referred to as the remnant.
The amplitudes $\mathcal{A}_{{\ell mn}}^{+/-}$ and phases $\phi_{{\ell mn}}^{+/-}$ characterise the excitation of each mode and are inferred from the data. The inclination of the black hole spin relative to the observer’s line of sight is given by $\iota$, and $\varphi$ denotes the azimuthal angle of the line of sight in the frame of the black hole, and without loss of generality is usually set to zero [65]. $S_{{\ell mn}}$ are the spin-weighted spherical harmonics [78] and $t_{0}$ is the reference start time. The oscillation frequencies and decay times are given by $\omega^{R}_{-2{\ell mn}}=2\pi f_{-2{\ell mn}}$ and $\omega^{I}_{-2{\ell mn}}=1/\tau_{-2{\ell mn}}$ respectively. The strain time series measured at the gravitational wave detector is a linear combination of both the plus ($h_{+}$) and cross ($h_{\times}$) polarizations [71], defined by
@@
where $F_{+}=F_{+}(\alpha,\delta,\psi)$ and $F_{\times}=F_{\times}(\alpha,\delta,\psi)$ are the detector response functions characterized by the two polarizations [79]. Both of these functions depend on the parameters $(\alpha,\delta,\psi)$ which are the sky localization parameters: right ascension, declination and polarization angle respectively. The sky location is fixed to coincide with the maximum likelihood value determined from the full Inspiral-Merger-Ringdown (IMR) analysis [80]. These detector values can be found for any ringdown event in the corresponding LVK configuration file [81, 82].
@@
$\omega_{-2{\ell mn}}|_{Kerr}\longrightarrow\omega_{0{\ell mn}}|_{EW1}.$ (40)
@@
We conclude this subsection with a reminder of some additional caveats to the analysis. The restriction of Kerr221 or EW221 to just a superposition of the fundamental QNM and its first overtone is an assumption. Although contributions from the overtones close to the peak have not been shown to correspond to any physical vibrational frequencies of the underlying geometry [15], their usage here will be to facilitate fitting the waveform close to the signal peak for sufficient accuracy due to the low signal-to-noise ratio (SNR) in many ringdown signals [79] observed by LVK. The inclusion of the overtone enables the analysis to be pushed to an earlier start time, capturing more of the power which is present near the peak of the signal. This is possible since the overtones are necessarily more damped than their fundamental counterparts, and can serve to effectively remove non-linearities present near the waveform peak, albeit at the risk of potentially removing interesting contributions due to the merger and thus ‘overfitting’ the data [15]. A second point to note is that in ringdown-only analyses, the precise value of the ringdown start time $t_{0}$ can have a non-negligible effect on the analysis and usually requires marginalization over its uncertainty. However, since our analysis will involve a fixed set of modes, where the primary goal is simply to consider whether the remnant is magnetized or not, the impact of the start time should be much smaller and, as in [79], we will fix the start time to reduce the computational cost.
@@
Figure 5: Posterior distributions for the remnant mass $M_{f}$, spin $\tilde{a}_{f}$, and magnetic field $\tilde{B}_{f}$ resulting from the analysis of four gravitational wave signals of binary mergers: GW150914, GW190521_074359, GW190727_060333 and GW200224_222234. The upper and lower limits on the estimated parameters lie within the $90\%$ credible interval.
@@ 4.2 Time-domain likelihood
[…]
As in [65], the likelihood will be assumed to follow a zero-mean Gaussian distribution with a standard deviation equal to the numerical uncertainty along with uniform priors on the coefficients of the template. The underlying stochastic sampling used to calculate all the posterior distributions in pyRing is performed using a nested sampling algorithm [83] implemented through the python package CPNest [84]. Once these posterior samples have been generated it is then possible to generate the marginalized posteriors for any subset of the parameters by simply selecting the corresponding samples - this property of marginalized distributions can be used to visualize the output of these samples by constructing corner plots, which show the marginalized one-dimensional and two-dimensional posterior distributions for each of the parameters.

### v3

@@ 3.1 Numerical QNM results
The equation can be solved using a root finding algorithm beginning with some initial guess $\omega^{0}_{{\ell mn}}$. To ensure convergence of the mode $\omega_{\ell mn}$ we will use the technique outlined in [Jaramillo:2020tuu, Siqueira:2022tbc, bohra2023gravitational] to fine-tune the required number of terms $N$ needed in the continued fraction. In each step we begin with the starting value of $N$ and determine $\omega_{\ell mn}$, increasing $N$ by $\delta N$, until the relative difference between the last two adjacent values meets a specified precision threshold. That is, for a chosen set of parameters $\lambda$, if $\omega_{\ell mn}(N)$ denotes the QNM frequency determined with an $N$-term continued fraction, the iteration will stop when
@@
Earlier calculations in the literature for Kerr [Leaver:1985analyticQNMKerr, Berti:2009kk] and Ernst [Konoplya:2007yy, Becar:2022wcj] black holes were used to validate our implementation of the continued fraction method. The non-Hermitian and thus inherently dissipative nature of the QNM (quasi)spectrum calculation requires the use of high machine precision throughout intermediate steps to avoid rounding errors, and in this work we have set the convergence criterion to $\epsilon=-7$.
[…]
A notable feature is that increasing $\tilde{B}$ leads to a reduction in the imaginary part of the QNM frequency in all cases. This reflects the tendency of the pure Melvin geometry to define box-like boundary conditions and exhibit normal as opposed to quasinormal modes. Another notable feature is the symmetry between positive and negative branches of modes. An analogous symmetry was observed by Leaver in [Leaver:1985analyticQNMKerr] in the case of the Kerr geometry, and can be understood in a similar way. In particular, if $\omega_{\ell mn}$ is a QNM frequency with azimuthal index $m$, then $\omega_{\ell-mn}=-\omega_{\ell mn}^{*}$. Similarly, the linearized EW equations contain combinations of $m^{2}\tilde{B}^{2}$, and $m\tilde{a}\tilde{B}^{2}$ meaning that if $m\rightarrow-m$ then $m^{2}\tilde{B}^{2}$ should remain invariant while $m\tilde{a}\tilde{B}^{2}\rightarrow-m\tilde{a}\tilde{B}^{2}$. These symmetries are apparent on computing the positive and negative branch of the quasinormal mode frequencies, as shown in Fig. 3.
@@ 3.2 Numerical QNM interpolation
where $\omega_{\ell 0n}$ is the Schwarzchild QNM for a given  $\ell\text{ and }n$ as listed in Table 1. We compute the coefficients $a_{\ell mn}$, $b_{\ell mn}$, and $c_{\ell mn}$ by taking numerical derivatives of the data obtained in the preceding section. In particular, $a_{\ell mn}$ describe the first derivative of the Kerr spectrum in the linearized regime of $\tilde{a}$, $b_{\ell mn}$ describe the first derivative of the Ernst spectrum in the linearized regime of $\tilde{B}^{2}$, and $c_{\ell mn}$ result from the numerical second derivative with respect to both $\tilde{a}$ and $\tilde{B}^{2}$. The results for these coefficients are tabulated in Tables 2 and 3, with those restricted to linearized Kerr agreeing with [Cardoso:2019mqo]. A comparison between the modes obtained directly from the Ernst geometry at ${\cal O}(\tilde{B}^{2})$ and the coefficients $b_{lmn}$ is shown in Fig. 4.
@@
1 1-1 0 -0.097660 -0.000336 0.268150 0.318134
11 -0.097660 0.000336 0.268150 -0.318132
2 2-2 0 -0.096759 -0.000143 0.432264 0.535567
1-1 -0.096759 -0.000072 0.108133 0.132246
11 -0.096759 0.000072 0.108133 -0.132246
22 -0.096759 0.000143 0.432264 -0.535561
2 2-2 1 -0.295604 -0.004641 0.933283 1.131990
1-1 -0.295604 -0.002321 0.233458 0.322617
11 -0.295604 0.002321 0.233458 -0.322418
22 -0.295604 0.004642 0.933283 -1.132600
3 3-3 0 -0.096500 -0.000078 0.512260 0.653273
2-2 -0.096500 -0.000052 0.227786 0.263341
1-1 -0.096500 -0.000026 0.056964 0.080038
11 -0.096500 0.000026 0.056964 -0.080027
22 -0.096500 0.000052 0.227786 -0.263310
33 -0.096500 0.000078 0.512260 -0.653221
4 4-4 0 -0.096392 -0.000048 0.558610 0.726651
3-3 -0.096392 -0.000036 0.314251 0.370947
2-2 -0.096392 -0.000024 0.139677 0.164431
1-1 -0.096392 -0.000012 0.034921 0.057358
11 -0.096392 0.000012 0.034921 -0.057357
22 -0.096392 0.000024 0.139677 -0.164431
33 -0.096392 0.000037 0.314251 -0.370944
44 -0.096392 0.000049 0.558610 -0.726644
Presently, there is some debate within the gravitational wave literature regarding which modes can be reliably extracted from the black hole ringdown. Indeed, it is clear that the process of ringdown following a physical binary merger is complex, and nonlinear effects may be as large as higher-order QNM contributions. In [Baibhav:2023clw], it is argued that even in the ideal situation when a binary black hole system collides head-on, only the modes $\{(2,2,0),(3,2,0),(4,2,0)\}$ significantly contribute to the strain. Reliable extraction of the first overtone $(2,2,1)$ would then depend on a high-precision determination of the longer-lived fundamental modes of subdominant multipoles. In the non-ideal case when the binary does not collide head-on, it becomes significantly more difficult to extract overtones and the dominant modes excited could differ from $\{(2,2,0),(3,2,0),(4,2,0),(2,2,1)\}$. For this reason we include the fit coefficients for the fundamental modes from $1\leq\ell\leq 4$ as well as the first overtone for $\ell=2$ to cover the range expected to be relevant in a range of ringdown analyses.
@@ 4 Ringdown template and comparison to LVK data
In this section we utilize our QNM results for the linearized EW geometry to build and test a ringdown template against the existing LVK data archive of binary black hole mergers. This serves as a first step in modifying the usual Kerr template to model the post-merger phase of a binary black hole coalescence in the presence of an external asymptotically uniform magnetic field. We model our approach on [Carullo:2021oxn], which considered the ringdown phase of black holes in the presence of a remnant $U(1)$ charge. However, we emphasize that this is a scalar analogue template and only precise for small $\tilde{B}$. Thus it cannot be used to constrain the magnetic field scale from data. Our exploratory approach will simply be to study $B$ as a nuisance-type parameter, and the template will be obtained by simply extrapolating the perturbative QNM spectrum.
To investigate the magnetized Ernst-Wild hypothesis in the current LVK data archive, we make use of pyRing [pyRing, Carullo:2019flw, Isi_2019], a python software package specifically designed for the estimation of black hole ringdown parameters. pyRing implements a Bayesian analysis approach, described within the LVK analysis pipeline in Refs. [LIGOScientific:2019hgc, LIGOScientific:2020tif, Isi:2021iql, Thrane_2019, Dreyer:2003bv, Gennari:2023gmx].
@@ 4.1 Constructing the ringdown template
To construct a waveform model for a black hole in the presence of an external magnetic field, we start from the standard Kerr template [Carullo:2021oxn, Berti:2005ys, Lim:2019xrb]:
@@
Here, $\tilde{\omega}_{s{\ell mn}}$ are the complex tensor (s=2s=-2) ringdown frequencies of the Kerr black hole determined by the dimensionless final spin $\tilde{a}_{f}$ and the detector frame final mass given by $M_{f}$ defined in [LIGOScientific:2016vlm]. $M_{f}$ is related to the source frame mass by a factor of $(1+z)$, and $D_{L}$ is the luminosity distance. The $f$ subscript on parameters in this section indicates that these values refer to the final black hole of the merger, commonly referred to as the remnant.
The amplitudes $\mathcal{A}_{{\ell mn}}^{+/-}$ and phases $\phi_{{\ell mn}}^{+/-}$ characterise the excitation of each mode and are inferred from the data. The inclination of the black hole spin relative to the observer’s line of sight is given by $\iota$, and $\varphi$ denotes the azimuthal angle of the line of sight in the frame of the black hole, and without loss of generality is usually set to zero [Carullo:2021oxn]. $S_{{\ell mn}}$ are the spin-weighted spherical harmonics [Berti:2014fga] and $t_{0}$ is the reference start time. The oscillation frequencies and decay times are given by $\omega^{R}_{-2{\ell mn}}=2\pi f_{-2{\ell mn}}$ and $\omega^{I}_{-2{\ell mn}}=1/\tau_{-2{\ell mn}}$ respectively. The strain time series measured at the gravitational wave detector is a linear combination of both the plus ($h_{+}$) and cross ($h_{\times}$) polarizations [Isi:2021iql], defined by
@@
where $F_{+}=F_{+}(\alpha,\delta,\psi)$ and $F_{\times}=F_{\times}(\alpha,\delta,\psi)$ are the detector response functions characterized by the two polarizations [mishra2023bounds]. Both of these functions depend on the parameters $(\alpha,\delta,\psi)$ which are the sky localization parameters: right ascension, declination and polarization angle respectively. The sky location is fixed to coincide with the maximum likelihood value determined from the full Inspiral-Merger-Ringdown (IMR) analysis [LIGOScientific:2021sio]. These detector values can be found for any ringdown event in the corresponding LVK configuration file [LIGO_data_2021_5172704, LIGO_data_2022_7007370].
@@
$\omega_{-2{\ell mn}}|_{\text{Kerr}}\longrightarrow\omega_{0{\ell mn}}|_{\text{EW1}}.$ (40)
@@
We conclude this subsection with a reminder of some additional caveats to the analysis. The restriction of Kerr221 or EW221 to just a superposition of the fundamental QNM and its first overtone is an assumption. Although contributions from the overtones close to the peak have not been shown to correspond to any physical vibrational frequencies of the underlying geometry [Baibhav:2023clw], their usage here will be to facilitate fitting the waveform close to the signal peak for sufficient accuracy due to the low signal-to-noise ratio (SNR) in many ringdown signals [mishra2023bounds] observed by LVK. The inclusion of the overtone enables the analysis to be pushed to an earlier start time, capturing more of the power which is present near the peak of the signal. This is possible since the overtones are necessarily more damped than their fundamental counterparts, and can serve to effectively remove non-linearities present near the waveform peak, albeit at the risk of potentially removing interesting contributions due to the merger and thus ‘overfitting’ the data [Baibhav:2023clw]. A second point to note is that in ringdown-only analyses, the precise value of the ringdown start time $t_{0}$ can have a non-negligible effect on the analysis and usually requires marginalization over its uncertainty. However, since our analysis will involve a fixed set of modes, where the primary goal is simply to consider whether the remnant is magnetized or not, the impact of the start time should be much smaller and, as in [mishra2023bounds], we will fix the start time to reduce the computational cost.
@@
Figure 5: Posterior distributions for the remnant mass $M_{f}$, spin $\tilde{a}_{f}$, and magnetic field $\tilde{B}_{f}$ resulting from the analysis of four gravitational wave signals of binary mergers from the LVK catalogue: GW150914, GW190521_074359, GW190727_060333 and GW200224_222234. The upper and lower limits on the estimated parameters bound the $90\%$ credible interval and are shown with the dashed purple lines.
@@ 4.2 Time-domain likelihood
The analysis goal is to construct the posterior distribution $p(\theta|d)$, where $\theta$ is the set of model parameters and $d$ is the data associated with the measurement. Bayes’ theorem determines $p(\theta|d)=\mathcal{L}(d|\theta)\,\pi(\theta)/\mathcal{Z}$ in terms of $\mathcal{L}(d|\theta)$, the likelihood function of the data given the parameters $\theta$, while $\pi(\theta)$ is the prior distribution for $\theta$ and $\mathcal{Z}\equiv\int d\theta\,\mathcal{L}(d|\theta)\,\pi(\theta)$ is the normalization factor, referred to as the ‘evidence’.
As in [Carullo:2021oxn], the likelihood will be assumed to follow a zero-mean Gaussian distribution with a standard deviation equal to the numerical uncertainty along with uniform priors on the coefficients of the template. The underlying stochastic sampling used to calculate all the posterior distributions in pyRing is performed using a nested sampling algorithm [Skilling:2004pqw] implemented through the python package CPNest [cpnest]. Each event sampler settings used in each event was 4096 live points, 4096 as a maximum number of Markov-Chain Montecarlo internal steps and a pool-size composed of 100 walkers. Increasing the live points and internal steps to 8192 did not impact the final results obtained. When the estimated precision on the logarithm of the evidence is within 0.1, the nested sampler stops. Once these posterior samples have been generated it is then possible to generate the marginalized posteriors for any subset of the parameters by simply selecting the corresponding samples - this property of marginalized distributions can be used to visualize the output of these samples by constructing corner plots, which show the marginalized one-dimensional and two-dimensional posterior distributions for each of the parameters.